# Unpacking Personal(?!) Health Informatics: An Investigation of Awareness, Understanding, & Leveraged Utility in India

### v1

@@ 3.1. Participants
The study recruited participants from across India, aiming to capture a diverse demographic range in age, gender, and professional background. To ensure a representative sample (especially those who can purchase and use even affordable smartphones and/or wearables), we focused on individuals from middle to upper-middle socioeconomic strata, categorized using the updated Kuppuswamy scale by Saleem et al. (Saleem and Jan, 2021). We employed a combination of purposive and snowball sampling methods to identify and engage participants. The sample included a variety of professions, such as IT professionals, healthcare workers, students, entrepreneurs, and retirees, reflecting the broad spectrum of urban Indian society. This approach enabled us to gather insights from participants with differing levels of familiarity and engagement with wearable health technologies. Prior to participation, all individuals were provided with a clear explanation of the study’s purpose, and informed consent was obtained. Demographic details for the surveys and interview participants are summarized in Table 1, offering a comprehensive overview of the sample’s composition.
Study Component | Total | Gender | Age | Education | |||||||
Male | Female | Prefer Not | Range | Mean | High | Bachelors | Advanced | ||||
to Say | School | Degrees | |||||||||
PHI awareness survey | 87 | 50 | 36 | 1 | 18-83 | 32 | 6 | 31 | 50 | ||
34 | 20 | 14 | 0 | 18-62 | 32 | 1 | 16 | 17 | ||
interviews | 22 | 16 | 8 | 0 | 18-62 | 28 | 5 | 6 | 11 | ||
co-design workshops | 12 | 6 | 6 | 0 | 27-40 | 33 | 0 | 8 | 4 | ||
114 | 44 | 70 | 0 | 19-75 | 34 | 3 | 57 | 54 |
[The table listing participant statistics from various study components]Table 1 lists the ranges and mean of age, gender, and educational background for the three components of the study.
@@ 3.2.1. Survey Design
The initial PHI awareness survey was administered using SurveyMonkey and targeted a diverse demographic of urban Indian participants. The survey included closed-ended and open-ended questions designed to capture a broad spectrum of wearable health technology use data. The survey comprised a total of 29 questions, including 8 questions on demographic information and 8 open-ended questions (6 for users and 2 for non-users). The closed-ended questions focused on collecting demographic information (age, gender, occupation), the type of wearable devices used for PHI, frequency of use, and specific health metrics tracked (such as steps, heart rate, and sleep patterns). The open-ended questions focused on aspects of trust and usage, such as (1)Please explain briefly if you have ever changed the health wearable you use and why. (2) What do you understand about the health data/analysis from your wearable? (3) Do you trust the accuracy of the wearable health data? (4) Has the history from your health wearable ever helped you with doctor visits? The open-ended questions of the survey allowed participants to elaborate on their experiences with wearable devices as part of their PHI, providing deeper insights into their motivations, the challenges they faced, and their overall perceptions of the utility of these devices in their daily lives. These responses were crucial for identifying recurring themes and patterns in the qualitative analysis. The survey remained open until saturation was observed in the responses.
@@
The PHI awareness survey had a total of 99 responses, of which 87 were complete and were used for further analysis. The demographics had a mean age of 32 and a maximum age of 83, which included 50 males, 36 females, and 1 participant who preferred not to disclose their gender. Among the complete responders, there were 55 users and 32 non-users. The majority (85%) of the survey participants were between 18 and 44 years old. The diversity of participants in age, gender, profession, and education provided a robust dataset for analyzing broader adoption/use/behavior patterns and informed the qualitative exploration in the subsequent interviews.
The concluding survey on health-sharing features was designed to assess both awareness and utilization of these features among Android and iOS users. It comprised 26 items in total: five demographic questions followed by seven feature-specific questions tailored to each subgroup (Apple users, Apple non-users, Android users, Android non-users). Consequently, participants who actively used the relevant health application responded to six closed-ended questions and one open-ended question, while non-users completed five closed-ended questions and two open-ended questions. The closed-ended items measured participants’ familiarity with the sharing function within their health-tracking app, their understanding of its capabilities, and—when applicable—the frequency of their usage. Open-ended questions invited respondents to elaborate on the reasoning behind their selected responses. The survey was closed upon reaching saturation of responses.
The health sharing feature survey was completed by 114 participants, with a mean age of 34 and a maximum age of 75, which included 44 males and 70 females. Among the responders, only 20 were users of the health sharing feature, 54 were non-users, and 40 did not even know such a feature existed. This final survey helped reinforce the findings of the interviews and co-design workshops on the lack of such sharing mechanisms, especially with the Android users (who are % of users in India).
ID | Age | Gender | Education | Wearable Usage | Was Gifted | Abandonment | Health Analytics |
P1 | 19 | Male | High School | Health | Yes | Still Using | Advanced |
P2 | 25 | Female | Advanced Degree | Non-User | N/A | N/A | N/A |
P3 | 33 | Female | Advanced Degree | Health | Yes | Still Using | Advanced |
P4 | 33 | Female | Bachelors | Health/Utility | Yes | Stopped Using | Advanced |
P5 | 29 | Female | Bachelors | Utility | No | Still Using | N/A |
P6 | 20 | Male | High School | Fashion | No | Stopped Using | N/A |
P7 | 32 | Male | Doctorate | Health/Utility | No | Still Using | Basic |
P8 | 26 | Male | Advanced Degree | Health | No | Still Using | Basic |
P9 | 29 | Female | Bachelors | Health/Utility | Yes | Stopped Using | Advanced |
P10 | 62 | Male | Bachelors | Health/Utility | Yes | Still Using | Basic |
P11 | 34 | Male | Bachelors | Health/Utility | No | Stopped Using | Advanced |
P12 | 20 | Male | High School | Fashion | No | Stopped Using | N/A |
P13 | 29 | Male | Advanced Degree | Health/Utility | Yes | Still Using | Advanced |
P14 | 20 | Male | High School | Non-User | N/A | N/A | N/A |
P15 | 25 | Female | Advanced Degree | Fashion/Health | Yes | Still Using* | Basic |
P16 | 26 | Female | Advanced Degree | Health/Fashion | No | Still Using | Basic |
P17 | 26 | Male | Advanced Degree | Health/Utility | Yes | Still Using | Basic |
P18 | 18 | Male | High School | Health/Utility | No | Still Using | Advanced |
P19 | 36 | Male | Advanced Degree | Health/Utility | No | Stopped Using** | Advanced |
P20 | 29 | Male | Advanced Degree | Non-User | N/A | N/A | N/A |
P21 | 27 | Male | Bachelors | Health/Utility | Yes | Still Using | Basic |
P22 | 27 | Female | Advanced Degree | Health/Utility | Yes | Stopped Using | Basic |
[The table listing participant demographics]Table 2 lists details of age, gender, education, usage/non-usage, level of health analytics, and Usage Abandonment for each participant by the uniquely assigned participant ID from the semi-structured Interviews.
@@ 3.2.2. Interview Design
The secondary component consisted of semi-structured interviews with 22 participants, all of whom were selected from the survey pool. The interviews were conducted based on the availability of the participants and were conducted until response saturation was reached on the emerging themes. Participants were chosen to reflect on various experiences with wearable technology for their PHI. The interview group included 16 males and 8 females, with a mean age of 28, from diverse educational backgrounds, including those with up to high school education, bachelor’s degrees, and advanced degrees. The detailed demographics are presented in Table 2. The interview group included twelve participants who actively tracked their health using wearables; the rest had never used a smart wearable for PHI or stopped using their wearable devices. This diversity in the interview sample allowed for a comprehensive exploration of the different attitudes and behaviors toward wearable health technologies for PHI.
The interview questions were designed to delve deeper into themes that emerged from the survey data, such as the influence of cultural practices on wearable adoption, perceptions of data accuracy, and usability challenges associated with wearables. The interviews also explored participants’ broader health management practices and how wearables for PHI fit within this ”ecology of tracking.” This approach allowed the researchers to capture rich, contextual data that provided insights beyond what could be obtained from the survey alone.
ID | Age | Gender | Education | No. of Children |
@@
[The table listing participant demographics of card sorting activity.]Table 3 lists details of age, gender, education, and number of children for each participant by the uniquely assigned participant ID from the co-design workshops.Advanced degree includes graduate degrees (like M.A., L.L.M., M.Sc., MS, MTech., M.Eng., MBA, CA, MD, MPhil, PhD, DPhil, etc.)
[Sample Card Sorting Activity Screens.]The screenshot shows the interface of the card sorting activity used by couples to group cards comprising of different health data (such as heart rate, SpO2, menstrual cycles, etc.) and aspects of sharing (such as control, privacy, and storage, etc.) to be sorted into categories of various stakeholder groups (such as parents, siblings, friends, doctors, etc.)
3.2.3. Co-design Workshops
One of the critical findings from the surveys and interviews was the participants’ desire to share health informatics with different stakeholders and the request for reliable tools to understand health information. This led us to consider ideating the interaction for sharing data with stakeholders of an individual’s PHI, such as Spouses, Parents/Adult Children, and Doctors, through the co-design workshops. To understand how, what, why, and when they would like to share, we conducted interviews with Spousal Pairs/Couples (6), Older Parent/Adult Children pairs (4), and doctors (4 from 3 specialties). To further identify the sharing preferences, we conducted a card sorting activity using kartSort (just using the couples as older participants struggled with the activity in the digital format and in-person activity was unfeasible as they lived across India) where they moved around different data/information cards to different stakeholders as categories. A sample screen from the card sorting activity is presented in Figure 1. We recruited six couples to reach saturation in understanding the variation of preferences among couples with and without a child (or children). The detailed demographics of the card sorting pairs are presented in Table 3.
[Sample Rapid Prototyping Screens.]The sample screenshot of the Figma prototype shows different screen flows in the prototype for selecting sharing preferences and approving requests for sharing among family members.
@@ 3.2.4. Rapid Prototyping and Evaluation
After the co-design workshops, we developed a medium fidelity prototype using Figma and evaluated its interaction with sharing preferences among couples and family members. This process helped us understand the ideal sharing preferences of individuals with various stakeholders involved. Figure 2 gives a sample screenshot of the prototype. The figma prototype was evaluated informally by sharing it with the participants and eliciting a qualitative evaluation. The unanimous positive takeaway from the participants’ feedback included its simplicity, modularity, and ease of control. Apart from these, there were individual features that the participant appreciated and provided further recommendations, which are discussed in the findings of this work. It was also noteworthy that participants from both Interviews and workshops generally lacked awareness of the health data sharing functionalities integrated into both Android and Apple smartphone applications, which indicated a need for another survey to gauge awareness, usage/non-usage of the same. The details of those observations are also presented in the findings, and the demographics were presented in 1.
Themes | Sub-Themes | Main Codes |
Collective health management | Use as a care device | Tracking heart rate & step count |
through sharing practices | Using because it was a gift | |
Track parents/partners health metrics | ||
Use as a health indicator | Use to monitor stress | |
Use to check critical conditions | ||
Step tracking as primary use | ||
Sleep tracking awareness | ||
Specific health tracking need | ||
(e.g., menstrual cycle) | ||
Digital PHI as a motivational tool | Perception on Wearable | Handy tool for tracking and utility |
for proactive health management | Useful motivational tool | |
Perception affecting degree of use | ||
Perceived low utility | ||
Trust in Data & Technology | Conditional trust in certain metrics | |
Acceptance of minor inaccuracies | ||
Useful as an approximate indicator | ||
Validation through comparison | ||
Low medical literacy on wearable data | ||
Navigating health information | Technical and Logistical Issues | Analytics not easy to understand |
in a fragmented digital ecosystem | Device management issues | |
Device incompatibilities | ||
Device change issues | ||
Buying a wearable is an investment | ||
Lifestyle and Physical Barriers | Lack of preferred modality of wearable | |
Rashes from the strap | ||
Minimal impact on doctor visits | ||
Physical discomfort on consistent use | ||
Preference for traditional methods | Trust on Body feeling | Will go to the doctor |
and perception of non-necessity | if I feel something wrong | |
Trust in health lifestyle/habits | ||
Perceptive lack of current need | ||
Reluctance to trust something new | ||
Trust in Doctors and | Will choose doctor for critical issues | |
digital PHI overwhelming | Do not know enough, doctors reliable | |
Wearable use overwhelming | ||
reactive care is enough | ||
Difference between body feeling | ||
and wearable metrics |
[Thematic Analysis table.]Table 4 lists the main themes, sub-themes, and the main codes obtained from the reflexive thematic analysis from the interview transcripts and the field notes taken during the interviews.
@@ 3.3. Data Analysis
We employed a combination of mixed methods to analyze the data. The survey responses were processed using descriptive statistics, where we calculated frequencies and percentages to assess patterns in the use and adoption of wearable health technologies for PHI, along with participants’ trust levels and perceived usefulness of these devices. To explore associations between demographic characteristics (such as age and gender) and usage patterns, we conducted cross-tabulations to identify significant relationships. We followed Braun and Clarke’s (Braun and Clarke, 2006) thematic analysis approach to comprehensively examine the interview data, open-ended survey responses, and field notes for the qualitative analysis. The analysis proceeded in several stages. Initially, the interview transcripts and field notes were reviewed multiple times to immerse the researchers in the data and identify recurring patterns. Subsequently, preliminary codes were developed to capture key concepts and insights. These codes were refined and organized into broader, more comprehensive themes. From this iterative process, four principal themes emerged: (1) Collective health management through sharing practices; (2) navigating health information in a fragmented digital ecosystem; (3) digital PHI as a motivational tool for proactive health management; and (4) preference for traditional methods and perception of non-necessity. These themes highlighted how PHI is viewed and used within the context of personal health management among urban Indian participants. The thematic coding framework is detailed in Table 4, providing a structured overview of the identified themes and their subcategories.
@@ 3.4. Data Triangulation
To ensure the robustness and credibility of our findings, we employed data triangulation(Carter et al., 2014) by integrating data from surveys, semi-structured interviews, co-design workshops, and existing literature. This method enabled cross-verification and identification of patterns and discrepancies across data types, enhancing the validity of the results. For instance, survey and interview data revealed a lack of trust in wearable technologies due to perceived inaccuracies in advanced metrics. This notion was corroborated by existing literature that links health literacy and trust to digital health technologies (Bagalkot et al., 2020; Kreps, 2017). Our analysis of barriers to wearable adoption, such as a preference for traditional healthcare and the need for culturally tailored information, aligns with broader studies emphasizing context-specific approaches in non-Western settings (Ugargol and Bailey, 2018; Volpato et al., 2021). Further, we observed that participants often viewed wearables as supplementary rather than essential health tools for their PHI, reflecting a prevalent reactive approach to health management as shown by prior literature (Lee et al., 2022; Kang and Exworthy, 2022). By integrating multiple data sources, our study offers a comprehensive understanding of the adoption, use, and perceptions of wearable health technologies among urban Indian users, thereby enhancing the reliability and depth of our findings.

### v2

@@ 3.1. Participants
The study recruited participants from across India, aiming to capture a diverse demographic range in age, gender, and professional background. To ensure a representative sample (especially those who can purchase and use at least affordable smartphones and/or wearables), we focused on individuals from middle to upper-middle socioeconomic strata, categorized using the updated Kuppuswamy scale (the latest peer-reviewed SES scale widely used for India) by Saleem et al. (Saleem and Jan, 2021). We employed a combination of purposive and snowball sampling methods to identify and engage participants. The sample comprised a diverse range of professions, including IT professionals, healthcare workers, students, entrepreneurs, and retirees, reflecting the broad spectrum of urban Indian society. This approach enabled us to gather insights from participants with differing levels of familiarity and engagement with wearable health technologies. Prior to participation, all individuals were provided with a clear explanation of the study’s purpose, and informed consent was obtained. Demographic details for the surveys and interview participants are summarized in Table 1, offering a comprehensive overview of the sample’s composition.
Study Component Total Gender Age Education Male Female Prefer Not Range Mean High Bachelors Advanced to Say School Degrees Survey (RQ1) 87 50 36 1 18-83 32 6 31 50 open-ended questions (RQ1) 34 20 14 0 18-62 32 1 16 17 Interviews (RQ1) 22 16 8 0 18-62 28 5 6 11 co-design workshops (RQ2) 12 6 6 0 27-40 33 0 8 4 Survey (RQ2) 114 44 70 0 19-75 34 3 57 54
@@ 3.2.1. Survey Design (RQ1)
The initial survey (examining RQ1) was administered using SurveyMonkey and targeted a diverse demographic of urban Indian participants. The survey included both closed-ended and open-ended questions, designed to capture a broad spectrum of data on wearable health technology use. The survey comprised a total of 29 questions, including 8 questions on demographic information and 8 open-ended questions (6 for users and 2 for non-users). The closed-ended questions focused on collecting demographic information (age, gender, occupation), the type of wearable devices used for PHI, frequency of use, and specific health metrics tracked (such as steps, heart rate, and sleep patterns). The open-ended questions focused on aspects of trust and usage, such as (1)Please explain briefly if you have ever changed the health wearable you use and why. (2) What do you understand about the health data/analysis from your wearable? (3) Do you trust the accuracy of the wearable health data? (4) Has the history from your health wearable ever helped you with doctor visits? The detailed survey questionnaire is presented in Appendix B. The open-ended questions of the survey allowed participants to elaborate on their experiences with wearable devices as part of their PHI, providing deeper insights into their motivations, the challenges they faced, and their overall perceptions of the utility of these devices in their daily lives for proactive care. These responses were crucial for identifying recurring themes and patterns in the qualitative analysis. The survey remained open until saturation was observed in the responses.
@@
The initial survey (RQ1) had a total of 99 responses, of which 87 were complete and were used for further analysis. The demographics consisted of a mean age of 32 and a maximum age of 83, with 50 males, 36 females, and 1 participant who preferred not to disclose their gender. Among the complete responders, there were 55 users and 32 non-users. The majority (85%) of the survey participants were between 18 and 44 years old. The diversity of participants, in terms of age, gender, profession, and education, provided a robust dataset for analyzing broader adoption/use/behavior patterns, and informed the qualitative exploration in the subsequent interviews.
ID Age Gender Education Wearable Usage Was Gifted Abandonment Health Analytics P1 19 Male High School Health Yes Still Using Advanced P2 25 Female Advanced Degree Non-User N/A N/A N/A P3 33 Female Advanced Degree Health Yes Still Using Advanced P4 33 Female Bachelors Health/Utility Yes Stopped Using Advanced P5 29 Female Bachelors Utility No Still Using N/A P6 20 Male High School Fashion No Stopped Using N/A P7 32 Male Doctorate Health/Utility No Still Using Basic P8 26 Male Advanced Degree Health No Still Using Basic P9 29 Female Bachelors Health/Utility Yes Stopped Using Advanced P10 62 Male Bachelors Health/Utility Yes Still Using Basic P11 34 Male Bachelors Health/Utility No Stopped Using Advanced P12 20 Male High School Fashion No Stopped Using N/A P13 29 Male Advanced Degree Health/Utility Yes Still Using Advanced P14 20 Male High School Non-User N/A N/A N/A P15 25 Female Advanced Degree Fashion/Health Yes Still Using* Basic P16 26 Female Advanced Degree Health/Fashion No Still Using Basic P17 26 Male Advanced Degree Health/Utility Yes Still Using Basic P18 18 Male High School Health/Utility No Still Using Advanced P19 36 Male Advanced Degree Health/Utility No Stopped Using** Advanced P20 29 Male Advanced Degree Non-User N/A N/A N/A P21 27 Male Bachelors Health/Utility Yes Still Using Basic P22 27 Female Advanced Degree Health/Utility Yes Stopped Using Basic
@@ 3.2.2. Interview Design
The secondary component consisted of semi-structured interviews with 22 participants, all of whom were selected from the survey pool. The interviews were conducted based on the availability of the participants and continued until response saturation was reached regarding the emerging themes. Participants were selected to reflect on their diverse experiences with wearable technology for their PHI. The interview group consisted of 16 males and 8 females, with a mean age of 28, representing diverse educational backgrounds, including those with up to high school education, bachelor’s degrees, and advanced degrees. The detailed demographics are presented in Table 2. The interview group included twelve participants who actively tracked their health using wearables; the rest had never used a smart wearable for PHI or had stopped using their wearable devices. This diversity in the interview sample allowed for a comprehensive exploration of the different attitudes and behaviors toward wearable health technologies for PHI.
The interview questions were designed to delve deeper into themes that emerged from the survey data, such as the influence of cultural practices on wearable adoption, perceptions of data accuracy, and usability challenges associated with wearables for proactive care. While the initial interview guide was limited to users, based on reviewing the survey responses, we reoriented the study design to provide a deeper examination of the non-user perspective. This provided us with vital information on the factors that prevented them from becoming users, some of which helped us understand how PHI can be better aligned. The interviews also explored participants’ broader health management practices and how PHI usage for proactive care fit within this ”ecology of tracking.” This approach enabled the researchers to capture rich, contextual data, providing insights that went beyond what could be obtained from the survey alone. The detailed interview guide is presented in Appendix C.
| ID | Age | Gender | Education | Child(ren) |
@@
3.2.3. Co-design Workshops and Stakeholder Interviews
One of the critical findings from the surveys and interviews was the participants’ desire to share health informatics with different stakeholders and the request for reliable tools to understand health information for proactive care. This led us to consider ideating the interaction for sharing data with stakeholders of an individual’s PHI. To understand how, what, why, and when they would like to share, we conducted interviews with Spousal Pairs/Couples (6) as part of the co-design workshops. As part of the workshops, we conducted a card sorting activity using kartSort, where participants moved different data/information cards to various categories for different stakeholders. A sample screen from the card sorting activity is presented in Figure 2. The different data/information cards could be duplicated, and the couple pairs could create new categories if needed for further granularity. The stakeholder interviews were conducted with older parent-adult child pairs (n = 4) (to avoid digital burden and unavailability for in-person workshops) and doctors (n = 4 from 3 specialties) (as they preferred interviews over workshops). Details of the stakeholder interview guide are presented in Appendix D. The workshops were conducted until saturation for understanding the variation of preferences among couples with and without a child (or children). The detailed demographics of the card sorting pairs are presented in Table 3.
@@ 3.2.4. Rapid Prototyping and Evaluation
After the co-design workshops, we developed a medium-fidelity prototype using Figma and evaluated its interaction with sharing preferences among couples and family members. This process helped us understand the ideal sharing preferences of individuals, involving various stakeholders. Figure 3 gives a sample screenshot of the prototype. The figma prototype was evaluated informally by sharing it with the participants and eliciting a qualitative evaluation. The unanimous positive takeaway from the participants’ feedback included its simplicity, modularity, and ease of control. Apart from these, the participant also appreciated individual features and provided further recommendations, which are discussed in the findings of this work. It was also noteworthy that participants from both Interviews and workshops generally lacked awareness of the health data sharing functionalities integrated into both Android and Apple smartphone applications, which indicated a need for another survey to gauge awareness and usage/non-usage of these functionalities. The details of those observations are also presented in the findings, and the demographics were presented in 1.
3.2.5. Survey Design (RQ2)
The concluding survey (examining RQ2) on perception, awareness, and utilization of health-sharing features was designed to assess these features among Android and iOS users, based on a critical finding on the willingness to share PHI data among stakeholders. It comprised 26 items in total: five demographic questions followed by seven feature-specific questions tailored to each of the four subgroups (Apple users, non-Apple users, Android users, and non-Android users). Consequently, participants who actively used the relevant health application responded to six closed-ended questions and one open-ended question, while non-users completed five closed-ended questions and two open-ended questions. The closed-ended items measured participants’ familiarity with the sharing function within their health-tracking app, their understanding of its capabilities, and the frequency of their usage (when applicable). Open-ended questions invited respondents to elaborate on the reasoning behind their selected responses. The detailed survey questionnaire is presented in Appendix E. The survey was closed once the desired number of responses was reached.
This survey (RQ2) was completed by 114 participants, with a mean age of 34 and a maximum age of 75, including 44 males and 70 females. Among the responders, only 20 were users of the health sharing feature, 54 were non-users, and 40 did not even know such a feature existed. This final survey helped reinforce the findings of the interviews and co-design workshops on the lack of such sharing mechanisms (or the lack of awareness of such features), especially among Android users (who account for % of users in India).
@@ 3.3. Data Analysis
We employed a combination of mixed methods to analyze the data. The survey responses were analyzed using descriptive statistics, which involved calculating frequencies and percentages to identify patterns in the use and adoption of wearable health technologies for PHI, as well as participants’ trust levels and perceived usefulness of these devices. To explore associations between demographic characteristics (such as age and gender) and usage patterns, we conducted cross-tabulations to identify significant relationships. We followed Braun and Clarke’s (Braun and Clarke, 2006) thematic analysis approach to comprehensively examine the interview data, open-ended survey responses, field notes, and the co-design session data for the qualitative analysis. The analysis proceeded in several stages. Initially, the interview transcripts and co-design field notes were reviewed multiple times to immerse the researchers in the data and identify recurring patterns. Subsequently, preliminary codes were developed to capture key concepts and insights. These codes were refined and organized into broader, more comprehensive themes. From this iterative process, two principal themes emerged: (1) Barriers to Proactivity and (2) Enablers of Proactivity: Behaviours & Triggers. These themes highlighted how PHI is viewed and used within the context of proactive care among urban Indian participants. The thematic coding framework is detailed in Table A in the appendix, providing a structured overview of the identified themes, subthemes/codes, along with their definitions, evidence (corresponding component of the study), and design implications.
@@ 3.4. Data Triangulation
To ensure the robustness and credibility of our findings, we employed data triangulation(Carter et al., 2014) by integrating data from surveys, semi-structured interviews, co-design workshops, and existing literature towards the recommendations and discussion. This method enabled cross-verification and identification of patterns and discrepancies across data types, enhancing the validity of the results. For instance, survey and interview data revealed a lack of trust in wearable technologies due to perceived inaccuracies in advanced metrics. This notion was corroborated by existing literature that links health literacy and trust to digital health technologies (Bagalkot et al., 2020; Kreps, 2017). Our analysis of barriers to wearable adoption, such as a preference for traditional healthcare and the need for culturally tailored information, aligns with broader studies emphasizing context-specific approaches in non-Western settings (Ugargol and Bailey, 2018; Volpato et al., 2021). Further, we observed that participants often viewed wearables as supplementary rather than essential health tools for their PHI, reflecting a prevalent reactive approach to health management as shown by prior literature (Lee et al., 2022; Kang and Exworthy, 2022). By integrating multiple data sources, our study offers a comprehensive understanding of the adoption, use, and perceptions of wearable health technologies among urban Indian users, thereby enhancing the reliability and depth of our findings for proactive care use.